# Comments, corrections, and related references welcomed, as always! ed August 22, 2025

### v2

@@ 6. Some results and examples for groups, comparing
Each of the above examples shows that the first implication of Lemma 6.3 is not reversible. I was unsure whether the second of those implications might be reversible, but Alexander Olshanskiy provided Example 6.8 below, showing that for every prime $p$ there is a group of order $p^{5}$ for which that reverse implication fails, and Example 6.9, showing that for $p=2,$ there are also two such examples of order $p^{3}.$ Example 6.10 and the following it paragraph give examples of order $p^{3}$ for all odd primes $p.$ The referee observes that all these are what are called “extraspecial p\!p-groups” – p\!p-groups $G$ whose center is isomorphic to $\mathbb{Z}_{p},$ and such that the quotient of $G$ by its center has exponent $p$ [13].
@@
By construction, $G$ is a quotient, hence a subquotient, of $H\times H;$ but $G$ is not a subquotient of $H,$ since $H$ has smaller order. Hence $G$ is not Rhodes-prime with respect to direct products.
@@
In $G,$ the subgroup $M$ described in the second paragraph of the example is central, hence normal, and has order $p,$ hence is minimal. To see that it is the only minimal normal subgroup, note that any $g\in G$ that is not in $M$ is the image of an element $(h_{0},h_{1})\in H\times H$ such that at least one of $h_{0}$ or $h_{1}$ involves a nonidentity power of $a$ or of $b.$ Assuming without loss of generality that $h_{0}$ involves a nonidentity power of $a,$ we find that the commutator of $(h_{0},h_{1})$ with $(b,I)$ is a nonidentity power of $(c,I).$ Hence in $G,$ the commutator of $g$ with the image of $(b,I)$ is a generator of $M.$ Thus, every normal subgroup of $G$ not contained in $M$ contains $M,$ so $G$ is indeed monolithic with monolith $M.$
@@
Then $Q_{8}$ is a subquotient of $D_{4}\times D_{4},$ and $D_{4}$ a subquotient of $Q_{8}\times Q_{8},$ but neither $Q_{8}$ nor $D_{4}$ is a subquotient of the other, hence neither group is Rhodes-prime with respect to direct products.
@@
Within $Q_{8}\times Q_{8},$ let $x=(i,1)$ and $y=(j,j).$ It is easy to check by looking at first coordinates, and then at second coordinates, that $x$ and $y$ satisfy the relations of (6.9), hence the group they generate is a homomorphic image of $A.$ That group admits a homomorphism onto $Q_{8},$ given by projection to the first component; but $x^{2}\,y^{2}=(1,-1)$ is in the kernel of that homomorphism, hence the group generated by $x$ and $y$ must have larger order than $Q_{8};$ hence it can’t be a proper homomorphic image of $A,$ so it must be isomorphic thereto.
Likewise, within $D_{4}\times D_{4},$ consider the subgroup generated by $x=(p,1)$ and $y=(q,p).$ As in the preceding paragraph, we verify that this group is a homomorphic image of $A,$ and that the projection onto the first component maps it surjectively to $D_{4}.$ However, $y^{2}=(e,p^{2})$ is a nonidentity element of the kernel of this homomorphism, so again the whole group is of larger order than $Q_{8},$ and so must be isomorphic to $A.$
@@
(We remark that $D_{4}$ is a semidirect product, $p\,\,|\,\,p^{4}=e\!>\,\rtimes<\!q\,\,|\,\,q^{2}=e\!>\!.$ However, $p\,\,|\,\,p^{4}=e\!>$ is not the least nontrivial normal subgroup of $D_{4}$ – that is a subgroup thereof – so the above non-Rhodes-primeness result does not contradict Proposition 6.6(ii).)
Here is one more example, answering a question raised by the referee:
@@
For any odd prime $p,$ the monolithic group $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}$ (where a generator of $\mathbb{Z}_{p}$ acts on $\mathbb{Z}_{p^{2}}$ by multiplication by 1+p,1\!+\!p, and the monolith is the subgroup $p\,\mathbb{Z}_{p^{2}}\rtimes\{e\})$ is isomorphic to a subquotient of $\mathbb{Z}_{p^{2}}\times H,$ for $H$ the group of upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ with diagonal $I,$ as in Example 6.8. Hence $G$ is not Rhodes-prime with respect to direct products.
@@
For $\mathbb{Z}_{p^{2}}$ we will use additive notation, writing $\mathbb{Z}_{p^{2}}=\{0,1,\dots,\,p^{2}{-}1\},$ while for $H$ we will use the multiplicative notation of Example 6.8, writing $H=\{a^{i}\,b^{j}\,c^{k}\ |\ 0\leq i,\,j,\,k<p\}$
Within $\mathbb{Z}_{p^{2}}\times H,$ take the subgroup of elements such that the $\mathbb{Z}_{p^{2}}\!-$coordinate, and the exponent of $a$ in the H\!H\!-coordinate, agree modulo $p;$ and divide this subgroup by the congruence equating the central elements (0,c)(0,\,c) and (p,I)(p,\,I) (both of order $p).$
The resulting subquotient is a group of order $p^{3}$ in which the image of $(1,\,b)\in\mathbb{Z}_{p^{2}}\times H$ generates a subgroup of order $p^{2},$ and the image of (0,a)(0,\,a) generates a subgroup of order $p,$ conjugation by which sends the generator of the former subgroup to its 1+p\!1{+}p-th power; so this subquotient is isomorphic to $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}.$ But $G$ is clearly not isomorphic to a subquotient of either $\mathbb{Z}_{p^{2}}$ or $H.$
@@
For $p$ a prime, what objects of the category of finite p\!p\!-groups are Rhodes-prime with respect to direct products? Are there any other than the cyclic p\!p\!-groups $\mathbb{Z}_{p^{n}}$?
I also haven’t thought hard about whether there are any non-Rhodes-prime finite monolithic groups that are not p\!p\!-groups.
@@
But $G$ is not monolithic, since it has at least the two minimal normal subgroups $\mathbb{Z}_{q_{0}}$ and $\mathbb{Z}_{q_{1}}.$ (These are its only minimal normal subgroups if $q_{0}\neq q_{1}.$ If $q_{0}=q_{1},$ on the other hand, then $\mathbb{Z}_{q_{0}}\times\mathbb{Z}_{q_{1}}$ is a 2\!2\!-dimensional $\mathbb{Z}_{q_{0}}\!$-vector space, and has $q_{0}\!+\!1$ one-dimensional subspaces, and if we give $\mathbb{Z}_{p}$ the same action on both of the original copies of $\mathbb{Z}_{q_{0}},$ then all $q_{0}\!+\!1$ of those subspaces are minimal normal subgroups of $G.)$
@@
Note that the modified Rhodes-primeness conditions are not, a priori, either stronger or weaker than the unmodified conditions:  The implication defining each of the modified conditions has a strengthened hypothesis, that $G$ be isomorphic to a subgroup of a group $G_{0}\times G_{1}$ or $G_{0}\rtimes G_{1},$ not merely to a subquotient; but it has a similarly strengthened conclusion. For finite groups, the modified conditions turned out to be weaker than or equivalent to the unmodified conditions, but for not-necessarily-finite groups, or various classes thereof, the relationships may be different.
@@
Understanding the two sorts of Rhodes-primeness to be defined for semigroups and monoids as we define them in Convention 6.1 for groups (they are so defined in [11]), and likewise for the condition of Tarski-primeness (as is the case in [9]), the following five implications are clear. Are some or all of them reversible?
@@
(6.10) $G$ is Rhodes-prime with respect to semidirect products as a finite semigroup
$\implies$
$G$ is Rhodes-prime with respect to semidirect products as a finite monoid
$\implies$
$G$ is Rhodes-prime with respect to semidirect products as a finite group.
(6.11) $G$ is Rhodes-prime with respect to direct products as a finite semigroup
$\implies$
$G$ is Rhodes-prime with respect to direct products as a finite monoid
$\implies$
$G$ is Rhodes-prime with respect to direct products as a finite group.
(6.12) $G$ is Tarski-prime as a finite semigroup
$\implies$
$G$ is Tarski-prime as a finite monoid, equivalently, as a finite group.
@@
One can see from the Jónsson-Tarski Theorem mentioned above that a finite group is Tarski-prime as a monoid if and only if it is Tarski-prime as a group, which is why (6.12) above shows fewer possibly distinct conditions than the other two collections. In the final sentence, (6.12) should actually be expanded to a 3\!3\!-case implication like the others.
Incidentally, because a semigroup or monoid is not in general isomorphic to its opposite, there are actually two versions of the concept of semidirect product for these objects, one based on “action on the right” and the other on “action on the left”. But since the opposite of every semigroup or monoid is still a semigroup or monoid, results about each of these constructions imply the corresponding results about the other. Cf. [11, p. 24, first sentence of next-to-last paragraph].
@@ 7. An observation on Rhodes-primeness with respect to direct
products in arbitrary varieties of algebras.
The concept of Rhodes-primeness with respect to direct products, as defined in Convention 6.1(iv) above, makes sense in any variety of algebras. (In contrast, in such a variety there’s no obvious version of the concept a semidirect product of objects, hence no obvious version of the concept of Rhodes-primeness with respect to semidirect products.) We end this note with a result on finite Rhodes-prime algebras in a general variety of algebras.
@@
Assuming (i), suppose we are given $Y_{0},$ $Y_{1}$ as in the hypothesis of (ii). Then by the assumption $\in\textup{Var}(\{Y_{0},\,Y_{1}\}), XX$ is a homomorphic image of the free algebra on $n$ generators in $\textup{Var}(\{Y_{0},\,Y_{1}\}),$ which can be constructed as a subalgebra of $Y_{0}^{\textup{card}(Y_{0})^{n}}\!\!\!\times Y_{1}^{\textup{card}(Y_{1})^{n}}\!\!.$ This makes $X$ a subquotient of a finite direct product of copies of $Y_{0}$ and $Y_{1}.$ Applying the Rhodes-primeness condition (i) inductively, we conclude that that $X$ is a subquotient of $Y_{0}$ or of $Y_{1},$ proving (ii).
@@
8. Acknowledgements.
I am grateful to Jonathan Farley for his group-email pointing out that the result of [14] answers one of the questions listed at [9, p. 263], which drew my attention to that list of questions, in particular the one now answered in §2 above; to Ralph Freese and Walter Taylor for information about what is known about Tarski-primeness; to Benjamin Steinberg, Stuart Margolis and Edmond Lee for information about what is known about the two Rhodes-primeness conditions; to Alexander Olshanskiy for Proposition 6.5, Examples 6.8 and 6.9, and some pointers on terminology, and to the referee for further information about known results and terminology.

### v3

@@ 6. Some results and examples for groups, comparing
Each of the above examples shows that the first implication of Lemma 6.3 is not reversible. I was unsure whether the second of those implications might be reversible, but Alexander Olshanskiy provided Example 6.8 below, showing that for every prime $p$ there is a group of order $p^{5}$ for which that reverse implication fails, and Example 6.9, showing that for $p=2,$ there are also two such examples of order $p^{3}.$ Example 6.10 and the paragraph following it give examples of order $p^{3}$ for all odd primes $p.$ The referee observes that all these are what are called “extraspecial p\!p-groups”, namely, p\!p-groups $G$ whose center is isomorphic to $\mathbb{Z}_{p},$ and such that the quotient of $G$ by its center has exponent $p$ [13].
@@
By construction, $G$ is a quotient, hence a subquotient, of $H\times H;$ but $G$ is not a subquotient of $H,$ since $H$ has smaller order. Hence $G$ is not Rhodes-prime with respect to direct products in the category of finite groups.
@@
In $G,$ the subgroup $M$ described in the second paragraph of the example is central, hence normal, and has order $p,$ hence is minimal. To see that it is the only minimal normal subgroup, note that any $g\in G$ that is not in $M$ is the image of an element $(h_{0},h_{1})\in H\times H$ such that at least one of $h_{0}$ or $h_{1}$ involves a nonidentity power of $a$ or of $b.$ Assuming without loss of generality that $h_{0}$ involves a nonidentity power of $a,$ we find that the commutator of $(h_{0},h_{1})$ with $(b,I)$ is a nonidentity power of $(c,I).$ Hence in $G,$ the commutator of $g$ with the image of $(b,I)$ is a generator of $M.$ Thus, every normal subgroup of $G$ not contained in $M$ contains $M,$ so as $M$ is simple, $G$ is indeed monolithic with monolith $M.$
@@
Then $Q_{8}$ is a subquotient of $D_{4}\times D_{4},$ and $D_{4}$ a subquotient of $Q_{8}\times Q_{8},$ but neither $Q_{8}$ nor $D_{4}$ is a subquotient of the other, hence neither group is Rhodes-prime with respect to direct products in the category of finite groups.
@@
Within $Q_{8}\times Q_{8},$ let $x=(i,1)$ and $y=(j,j).$ It is easy to check by looking at first coordinates, and then at second coordinates, that $x$ and $y$ satisfy the relations of (6.9), hence the group they generate is a homomorphic image of $A.$ That group admits a homomorphism onto $Q_{8},$ given by projection to the first component; but $x^{2}\,y^{2}=(1,-1)$ is in the kernel of that homomorphism, hence the group generated by $x$ and $y$ must have larger order than $Q_{8};$ hence it can’t be a proper homomorphic image of the 16\!16\!-element group  $A,$ so it must be isomorphic thereto.
Likewise, within $D_{4}\times D_{4},$ consider the subgroup generated by $x=(p,1)$ and $y=(q,p).$ As in the preceding paragraph, we verify that this group is a homomorphic image of $A,$ and that the projection onto the first component maps it surjectively to $D_{4}.$ However, $y^{2}=(e,p^{2})$ is a nonidentity element of the kernel of that homomorphism, so again the whole group is of larger order than $Q_{8},$ and so must be isomorphic to $A.$
@@
(We remark that $D_{4}$ is a semidirect product, $p\,\,|\,\,p^{4}=e\!>\,\rtimes<\!q\,\,|\,\,q^{2}=e\!>\!.$ However, $p\,\,|\,\,p^{4}=e\!>$ is not the least nontrivial normal subgroup $M$ of $D_{4}$ – that is a subgroup thereof – so the above non-Rhodes-primeness result does not contradict Proposition 6.6(ii).)
Here is one more example, which answers a question raised by the referee:
@@
For any odd prime $p,$ the monolithic group $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}$ (where a generator of $\mathbb{Z}_{p}$ acts on $\mathbb{Z}_{p^{2}}$ by multiplication by 1+p,1\!+\!p, and the monolith is the subgroup $p\,\mathbb{Z}_{p^{2}}\rtimes\{e\})$ is isomorphic to a subquotient of $\mathbb{Z}_{p^{2}}\times H$ where, as in Example 6.8, $H$ is the group of upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ with diagonal $I.$ Hence $G$ is not Rhodes-prime with respect to direct products in the category of finite groups.
@@
In $\mathbb{Z}_{p^{2}}$ we will use additive notation, writing $\mathbb{Z}_{p^{2}}=\{[0],[1],\dots,\,[p^{2}{-}1]\},$ while in $H$ we will use the multiplicative notation of Example 6.8, writing $H=\{a^{i}\,b^{j}\,c^{k}\ |\ 0\leq i,\,j,\,k<p\}$
Within $\mathbb{Z}_{p^{2}}\times H,$ take the subgroup of elements such that the $\mathbb{Z}_{p^{2}}\!-$coordinate, and the exponent of $a$ in the H\!H\!-coordinate, agree modulo $p;$ and divide this subgroup by the congruence equating the central elements ([0],c)([0],\,c) and ([p],I)([p],\,I) (both of order $p).$
The resulting subquotient is a group of order $p^{3}$ in which the image of $([1],\,b)\in\mathbb{Z}_{p^{2}}\times H$ generates a subgroup of order $p^{2},$ and the image of ([0],a)([0],\,a) generates a subgroup of order $p,$ conjugation by which sends the generator of the former subgroup to its 1+p\!1{+}p-th power; so this subquotient is isomorphic to $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}.$ But $G$ is clearly not isomorphic to a subquotient of either $\mathbb{Z}_{p^{2}}$ or $H.$
@@
For $p$ a prime, what finite p\!p\!-groups are Rhodes-prime with respect to direct products in the category of finite groups? Are there any other than the cyclic p\!p\!-groups $\mathbb{Z}_{p^{n}}$?
On the other hand, are there any finite monolithic groups not Rhodes-prime with respect to direct products that aren’t p\!p\!-groups? Yes. Modifying a construction suggested by the referee, let us consider the variant of Example 6.8 gotten by taking $p>2,$ and letting $H$ consist of all upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ in which $e_{11}$ and $e_{33}$ have coefficient $1,$ but the coefficient of $e_{22}$ can be any nonzero member of $\mathbb{Z}_{p};$ thus $|H|=(p{-}1)\,p^{3}.$ Obtain $G,$ as before, by dividing $H\times H$ by the normal subgroup of elements $(I+fe_{1,3},\,I-fe_{1,3})$ for $f\in\mathbb{Z}_{p};$ so $|G|=(p{-}1)^{2}\,p^{5}.$ The reader can verify that the image of the subgroup of elements $(I+fe_{1,3},\,I)$ is again a monolith. As before, though $G$ is a subquotient of $H\times H,$ it is too large to be a subquotient of $H,$ hence it is not Rhodes-prime with respect to direct products. (One can generalize this construction in a way that allows the monolith to have order $2$: Replace $\mathbb{Z}_{p}$ by any finite field of cardinality >2,>2, in particular allowing finite proper extensions of $\mathbb{Z}_{2},$ and at the end of the construction, divide the central subgroup of elements $(I+fe_{1,3},\,I+g\,e_{1,3})$ not just by the subgroup of elements $(I+fe_{1,3},\,I-fe_{1,3}),$ but by a large enough overgroup thereof to shrink it down to a cyclic group, so that it again becomes a monolith.)
@@
But $G$ is not monolithic, since it has at least the two minimal normal subgroups $\mathbb{Z}_{q_{0}}$ and $\mathbb{Z}_{q_{1}}.$ (These are its only minimal normal subgroups if $q_{0}\neq q_{1}.$ If $q_{0}=q_{1},$ on the other hand, then $\mathbb{Z}_{q_{0}}\times\mathbb{Z}_{q_{1}}$ is a 2\!2\!-dimensional $\mathbb{Z}_{q_{0}}\!$-vector space, and has $q_{0}\!+\!1$ one-dimensional subspaces; and if we give $\mathbb{Z}_{p}$ the same action on both of those copies of $\mathbb{Z}_{q_{0}},$ then all $q_{0}\!+\!1$ of those subspaces are minimal normal subgroups of $G.)$
@@
Note that the modified Rhodes-primeness conditions are not, a priori, either stronger or weaker than the unmodified conditions: The implication defining each of the modified conditions has a strengthened hypothesis, that $G$ be isomorphic to a subgroup of a group $G_{0}\times G_{1}$ or $G_{0}\rtimes G_{1},$ not merely to a subquotient; but it has a similarly strengthened conclusion. For finite groups, the modified Rhodes-primeness conditions turned out to be weaker than or equivalent to the unmodified conditions, but for not-necessarily-finite groups, or various classes thereof, the relationships may be different.
@@
Understanding the two sorts of Rhodes-primeness to be defined for semigroups and monoids as we define them in Convention 6.1 for groups (indeed, they are so defined for semigroups in [11]), and likewise for the condition of Tarski-primeness (which is defined for general algebras in [9]), the following five implications are clear. Are some or all of them reversible?
@@
(6.10) $G$ is Rhodes-prime with respect to semidirect products as a finite semigroup $\implies$ $G$ is Rhodes-prime with respect to semidirect products as a finite monoid $\implies$ $G$ is Rhodes-prime with respect to semidirect products as a finite group.
(6.11) $G$ is Rhodes-prime with respect to direct products as a finite semigroup $\implies$ $G$ is Rhodes-prime with respect to direct products as a finite monoid $\implies$ $G$ is Rhodes-prime with respect to direct products as a finite group.
(6.12) $G$ is Tarski-prime as a finite semigroup $\implies$ $G$ is Tarski-prime as a finite monoid, equivalently, as a finite group.
@@
One can see from the Jónsson-Tarski Theorem mentioned above that a finite group is Tarski-prime as a monoid if and only if it is Tarski-prime as a group, which is why (6.12) above shows fewer conditions than the other two displays. (But in the final sentence of Question 6.14 the analog of (6.12) could be expanded to a 3\!3\!-condition implication like the others.)
Incidentally, because a semigroup or monoid is not in general isomorphic to its opposite, there are actually two versions of the concept of semidirect product for these objects, one based on “action on the right” and the other on “action on the left”. But since the opposite of every semigroup or monoid is still a semigroup or monoid, general results about each of these constructions imply the corresponding results about the other. Cf. [11, p. 24, first sentence of next-to-last paragraph].
@@ 7. An observation on Rhodes-primeness with respect to direct
products in arbitrary varieties of algebras
The concept of Rhodes-primeness with respect to direct products, as defined in Convention 6.1(iv) above, makes sense in any variety of algebras. (In contrast, in a general variety there’s no obvious version of the concept of semidirect product of objects, hence no obvious version of the concept of Rhodes-primeness with respect to semidirect products.) We end this note with a result on finite algebras Rhodes-prime with respect to direct products in general varieties of algebras.
@@
Assuming (i), suppose we are given $Y_{0},$ $Y_{1}$ as in the hypothesis of (ii). Then taking $n$ such that $X$ is generated by $n$ elements, the assumption $X\in\textup{Var}(\{Y_{0},\,Y_{1}\})$ makes $X$ a homomorphic image of the free algebra on $n$ generators in $\textup{Var}(\{Y_{0},\,Y_{1}\}),$ which can be constructed as a subalgebra of $Y_{0}^{\textup{card}(Y_{0})^{n}}\!\!\!\times Y_{1}^{\textup{card}(Y_{1})^{n}}\!\!.$ This makes $X$ a subquotient of a finite direct product of copies of $Y_{0}$ and $Y_{1}.$ Applying the Rhodes-primeness condition (i) inductively, we conclude that that $X$ is a subquotient of $Y_{0}$ or of $Y_{1},$ proving (ii).
@@
In that same condition (iii), can the inclusion $\textup{Var}(X)\subseteq\textup{Var}(Y)$ be replaced by $\textup{Var}(X)=\textup{Var}(Y)$? Here is a version of an example provided by the referee showing that it cannot. Let $G$ be a nontrivial finite group, $V$ the variety of G\!G\!-sets, and $X$ a 2\!2\!-element set on which $G$ acts trivially. Thus, $X$ is a subquotient of a G\!G\!-set $Y$ if and only if $Y$ has more than one orbit. In particular, if $Y$ is a free G\!G\!-set on one generator, then $X$ is not a subquotient of $Y,$ but is a subquotient of $Y\times Y;$ so $X$ is not Rhodes-prime with respect to direct products. But turning to condition (iii), we see that $\textup{Var}(X)$ is the variety of G\!G\!-sets on which $G$ acts trivially, which is a minimal nontrivial element of the lattice of subvarieties of $V,$ hence join-prime; and we easily see that the “whenever” condition of (iii) with $\subseteq$ replaced by == also holds for $X;$ so that modified version of (iii) is not equivalent to (i).
8. Acknowledgements
I am grateful to Jonathan Farley for his group-email pointing out that the result of [14] answers one of the questions listed at [9, p. 263], which drew my attention to that list of questions, in particular the one now answered in §2 above; to Ralph Freese and Walter Taylor for information about what is known about Tarski-primeness; to Benjamin Steinberg, Stuart Margolis and Edmond Lee for information about what is known about the two Rhodes-primeness conditions; to Alexander Olshanskiy for Proposition 6.5, Examples 6.8 and 6.9, and some pointers on terminology, and to the referee for further information about known results and terminology, and some useful examples and suggestions.